OpenStudy (anonymous):

how do you convert 17.3 cm3 (cubed) into liters

OpenStudy (anonymous):

ok this is simple one... just remember this: when ever you have cubed units you need to add or subbstract 10^3 or 1000 to a number, let me give an example to make it clear: 1 L = 1 dm^3 1 mL = 1 cm^3 so 1,5 cm^3 = 1,5 * 10^-3 dm^-3 or 1,5 * 10^-3 L or 0,001 L or 0,001 dm^3 if you have any doubts just ask

OpenStudy (anonymous):

what i said is worth but i forgot to say about larger differences... let me give you an example ( you need to sum up values ): 1,5 cm^3 = 1,5 * 10^-3 dm3 = 1,5 * 10^-6 m3 and so on... so you sum or substract 1000 or 10^3 hope this helps further more ;)

OpenStudy (anonymous):

.0173 lt (liters)